OpenStudy (gabylovesyou):

The formula for degrees Celsius is: C = (F – 32), where F stands for degrees Fahrenheit. Part 1: Solve the formula for F. Show all steps. Part 2: Determine how many degrees Fahrenheit 6 degrees Celsius is. PLEASE HELP MEEE -_-

OpenStudy (gabylovesyou):

@dumbcow

OpenStudy (anonymous):

First of all, C = (F - 32) * 5/9. You left that last part out. To solve for F: C = (F - 32) * 5/9 C * 9/5 = F - 32 (C * 9/5) + 32 = F F = C * 9/5 + 32 Next part: C = 6 F = C * 9/5 + 32 F = 6 * 9/5 + 32 F = 54/5 + 32 F = 10.8 + 32 F = 42.8 degrees

OpenStudy (gabylovesyou):

whoops!! and ohh ok thanks :)

OpenStudy (anonymous):

welcome